# AmodalGen3D: Generative Amodal 3D Object Reconstruction from Sparse Unposed Views

### v1

@@ Abstract
Reconstructing 3D objects from a few unposed and partially occluded views is a common yet challenging problem in real-world scenarios, where many object surfaces are never directly observed. Traditional multi-view or inpainting-based approaches struggle under such conditions, often yielding incomplete or geometrically inconsistent reconstructions. We introduce AmodalGen3D, a generative framework for amodal 3D object reconstruction that infers complete, occlusion-free geometry and appearance from arbitrary sparse inputs. The model integrates 2D amodal completion priors with multi-view stereo geometry conditioning, supported by a View-Wise Cross Attention mechanism for sparse-view feature fusion and a Stereo-Conditioned Cross Attention module for unobserved structure inference. By jointly modeling visible and hidden regions, AmodalGen3D faithfully reconstructs 3D objects that are consistent with sparse-view constraints while plausibly hallucinating unseen parts. Experiments on both synthetic and real-world datasets demonstrate that AmodalGen3D achieves superior fidelity and completeness under occlusion-heavy sparse-view settings, addressing a pressing need for object-level 3D scene reconstruction in robotics, AR/VR, and embodied AI applications.
@@ 1 Introduction
Reconstructing 3D objects from a few unposed and partially occluded images is a fundamental but challenging task in computer vision. In realistic scenarios, objects are often captured from arbitrary viewpoints, with large portions hidden by surrounding structures or other objects. Such incomplete observations make the reconstruction problem inherently amodal: a system must infer the geometry and appearance of regions that are never directly observed. This ability is crucial for object-level 3D scene understanding and downstream applications such as robotic interaction, AR/VR content creation, and embodied AI perception.
Figure 1: Comparison of 3D object reconstruction from sparse, unposed, and occluded inputs. From left to right: input sparse views, results from (a) MVS-based reconstruction (e.g., VGGT [wang2025vggt]), (b) 2D amodal completion lifted to 3D (TRELLIS [xiang2025structured]), and (c) AmodalGen3D (ours). Our generative model faithfully reconstructs complete, occlusion-free 3D geometry consistent with sparse-view constraints.
Traditional multi-view stereo (MVS) and neural rendering approaches, such as NeRF or Gaussian Splatting, rely on dense and accurately calibrated views to estimate intact and consistent geometry. When only a few unposed observations are available, these methods produce incomplete or misaligned reconstructions, as shown in Fig. 1(a). On the other hand, 2D amodal completion techniques can hallucinate missing regions in image space, but their outputs are inconsistent across views, leading to unstable or implausible 3D geometry when combined, as illustrated in Fig. 1(b). The key challenge lies in jointly reasoning about both visible and occluded regions while maintaining geometric and appearance consistency under sparse, unposed constraints.
@@
To address this challenge, we introduce AmodalGen3D, a generative framework for amodal 3D object reconstruction from sparse, unposed views. Our approach treats reconstruction as a conditional generation process that integrates 2D amodal priors with explicit 3D geometric reasoning. Specifically, we first perform object-centric amodal completion on each view to recover both visible and occluded regions. The completed views are then encoded by a DINO-based feature extractor and aggregated through a View-Wise Cross Attention module that maintains spatial consistency across sparsely sampled viewpoints. To infer missing geometry hidden from occluders, we condition the generation on partial geometry estimated from an MVS model, represented as a point cloud. This information is incorporated through a Stereo-Conditioned Cross Attention module equipped with a geometry-guided gating mechanism, which modulates feature fusion based on 3D spatial cues. After processing the generated sparse structure through another pretrained Flow Transformer, the resulting structured latent representation is then decoded by specific decoders to produce a complete, occlusion-free 3D model that respects the sparse-view constraints while plausibly reconstructing the hidden regions. As shown in Fig. 1(c), our method produces coherent and complete 3D reconstructions even when large parts of the object are never observed.
@@
Extensive experiments on both synthetic and real-world datasets demonstrate that AmodalGen3D achieves superior reconstruction quality compared with existing sparse-view and amodal baselines. Our method improves geometric accuracy while maintaining realistic appearance, achieving consistent performance from single-view to multi-view inputs. These results highlight the potential of generative modeling for practical 3D object-level scene reconstruction under real-world occlusion and view sparsity.
@@
  • We introduce AmodalGen3D, a generative framework for amodal 3D object reconstruction that can recover complete geometry and appearance from sparse, unposed, and partially occluded inputs.
  • We propose the View-Wise Cross Attention that benefits feature aggregation from sparse views, and the novel Stereo-Conditioned Cross Attention for integrating partial MVS geometry into the generation process.
  • Comprehensive experiments validate that AmodalGen3D achieves superior reconstruction fidelity, completeness, and consistency compared with existing sparse-view and amodal reconstruction methods.
@@ 2 Related Work
Multi-View Stereo (MVS). Multi-view stereo (MVS) [FurukawaHernandez2015_MultiViewStereoTutorial, Wang2024_LearningBasedMVS_Survey, Yi2020PyramidMVSNet, Yan2020DenseHybridRMVSNet, wang2024dust3r, leroy2024grounding, zhang2024monst3r, yang2025fast3r, zhang2025flare, wang2025pi, keetha2025mapanything] has been studied for decades as a core technique for 3D reconstruction from multiple input views. Recent advances [Ren_2023_HPMMVS, Wu_2024_GoMVS, Santo_2024_MVCPS_NeuS] in dense correspondence learning have enhanced MVS accuracy and robustness by jointly reasoning about geometry and radiance. These methods achieve fine-grained alignment but remain limited when it comes to calibrated settings and fixed depth ranges. To improve scalability and avoid per-scene optimization, feed-forward extensions [Izquierdo_2025_MVSAnywhere, Cabon_2025_MUSt3R, Liu_2024_MVSGaussian, Dong_2024_DepthRangeFreeMVS] introduce transformer-based architectures and specific techniques for zero-shot multi-view reconstruction. These designs collectively enhance scalability and inference efficiency. A complementary line of work [Rojas_2025_HAMSt3R, Peng_2023_GenS, Rich_2024_DIVLoss] explores unposed and unstructured generalization in MVS.
Despite these advances, outputs from MVS methods are often incomplete or noisy (e.g., point clouds), which lack in semantic and appearance consistency for object-level understanding. In contrast, our approach leverages MVS as the geometric prior explicitly: by conditioning the generation model on partial stereo point clouds, AmodalGen3D infers complete, occlusion-free object structures from sparse and unposed inputs, bridging the gap between recent feed-forward MVS models and object-level reconstruction.
Figure 2: Overview of AmodalGen3D. Given sparse images $S$, visibility masks $M_{vis}$, and occlusion masks $M_{occ}$ indicating the occluded object $O$, AmodalGen3D first generates a sparse structure (stage 1) by aggregating multi-view information in Sec. 3.1 and infers the complete geometric structure from the partial stereo point cloud $\mathbf{P}_{O}$ in Sec. 3.2. Once the sparse structure is obtained, we employ a pretrained amodal SLAT Transformer, controlling texture generation with visibility masks $M_{vis}$ and occlusion masks $M_{occ}$, and then generate the structured latent. This structured latent is decoded into an occlusion-free 3D object with high-quality geometry and appearance.
@@
3D Generative Models. Recent advances in 3D generative models have opened new possibilities for structure-aware reconstruction. Methods leveraging Gaussian Splatting [kerbl20233d] and NeRF [mildenhall2021nerf] as the representation excel at novel view synthesis [ye2024no, truong2023sparf], but struggle to segregate object-level entities, hindering object-level 3D scene understanding. Diffusion-based generative models further enhance visual fidelity, controllability, and scalability by leveraging pretrained 2D diffusion priors [huang2024dreamcontrol, ding2024bidirectional, chen2024vp3d] or 3D-aware latent representations [xiang2025structured, long2024wonder3d, zhao2025hunyuan3d, lin2025kiss3dgen, siddiqui2024assetgen]. These methods typically rely on massive optimization iterations or decently posed input images for proper results. Some works attempt to mitigate missing regions using inpainting-augmented generation [chen2024comboverse, han2024reparo, zhou2024layout]. Other scene-level generative pipelines [chen2024single, liu2022towards, huang2025midi, meng2025scenegen, yao2025cast] synthesize entire 3D scenes from a single image, but they rely on carefully curated data and pose-calibrated views, which limits their effectiveness under real-world, unposed scenarios.
@@
In contrast to prior approaches, AmodalGen3D is a generative 3D reconstruction model tailored for sparse, unposed inputs. By fusing generative priors with partial geometric conditioning, it produces reconstructions that are both geometrically faithful and semantically coherent, remaining consistent with sparse-view constraints while plausibly inferring unseen object regions.
@@
Amodal Completion in 2D/3D. Amodal completion aims to infer the hidden or occluded parts of objects beyond visible evidence, which is fundamental for robust scene understanding and reconstruction. Early studies address this through explicit amodal segmentation and completion tasks [LiM_2016_ECCV, Qi_2019_CVPR, Ling_2020_NeurIPS, Mohan_2022_CVPR, Li_2023_ICCV_MUVA], focusing primarily on 2D perception and annotation-based benchmarks. These methods rely on supervised mask prediction or latent shape priors to hallucinate occluded object regions, but lack in strong generative fidelity and geometric consistency.
@@
Generative 2D inpainting models [Rombach_2022_CVPR, yu2023inpaint] have since become common solutions in image-level amodal completion, synthesizing realistic missing content based on surrounding context. Nowadays, researchers further enhance this capacity by introducing specialized amodal reasoning modules [ao2025open, lee2025tuningfreeamodalsegmentationocclusionfree, fan2025multi] or by fine-tuning diffusion models for occlusion-aware reconstruction [ozguroglu2024pix2gestalt, xu2024amodal, lu2025taco]. Extending beyond single-view imagery, multi-view inpainting models [barda2025instant3dit, weber2023nerfiller] attempt to maintain cross-view coherence but often assume calibrated cameras or dense view coverage. More geometry-aware approaches [Li_2022_ECCV, salimi2025geometry, feng2025objfiller, shi2025imfine, Zhan_2024_CVPR] leverage 3D representation priors for scene-level completion; however, they remain limited in handling object-level amodal reconstruction due to alignment errors and weak semantic reasoning. The most relevant work is Amodal3R [wu2025amodal3r], which trains a native 3D generation model from occluded objects, but it is constrained by data scale and limited occlusion reasoning ability.
@@
In sparse and unposed scenarios, 2D inpainting models can recover locally plausible structures yet introduce cross-view inconsistencies that manifest as geometric artifacts when reconstructed to 3D. Our method integrates these powerful 2D amodal priors into a 3D generative framework, enforcing global geometric coherence through the view-wise and stereo-conditioned attention mechanisms. Compared with single-image 3D amodal methods, AmodalGen3D achieves robust reconstruction under severe occlusion, benefiting from multi-view information aggregation and strong cross-view geometric reasoning, while maintaining the realism of image-level completion.
@@ 3 Method
For an arbitrary object $O$ in the given $K$ sparse views $S=\{s^{1},s^{2},...,s^{K}\}$ (typically 1-3 views), we can obtain the visibility masks of $O$ and the corresponding text prompt automatically or manually using vision foundation models, e.g., SAM [kirillov2023segment] and Florence 2 [xiao2024florence]. After that, we follow the occlusion analysis and inpainting strategies from [ao2025open] with minor modifications (using the above obtained visibility mask and prompt) to first perform 2D amodal completion across all views. Altogether, we can obtain visual clues of the desired object $O$ at all $K$ input views: the visibility masks $M_{vis}=\{m^{1}_{v},m^{2}_{v},...,m^{K}_{v}\}$ indicating the visible part of $O$; the inconsistent inpainted views of $O$: $I=\{i^{1},i^{2},...,i^{K}\}$; and the occlusion masks $M_{occ}=\{m^{1}_{o},m^{2}_{o},...,m^{K}_{o}\}$ representing the potential occlusion regions.
We illustrate the overall model structure in Fig. 2 using two views as an example. Since geometric completion plays an important role in amodal 3D reconstruction, our design prioritizes the Sparse Structure Transformer, i.e., stage 1. Given object-centric sparse views $I$ processed using 2D amodal completion methods, we apply the proposed View-Wise Cross Attention (Sec. 3.1) to aggregate multi-view features and fuse their latents. The fused representation is further refined through the Stereo-Conditioned Cross Attention (Sec. 3.2) for unobserved structure inference, guided by observed partial stereo geometry and a geometry-aware gating mechanism. The resulting Sparse Structure is decoded and then passed to the pretrained Structure Latent Transformer from [wu2025amodal3r] for Structured Latent generation. Implementation details of the training/inference proxy and the data engine used to construct the object-centric synthetic dataset for model training are described in Sec. 3.3.
Figure 3: The TRELLIS model shows a strong bias and can result in artifact accumulation when conditioned with multiple views.

### v2

# https://colezwhy.github.io/gena3d/

@@ Abstract
Generating complete 3D objects under partial occlusions (i.e., amodal scenarios) is a practically important yet challenging problem, as large portions of object geometry are unobserved in real-world scenarios. Existing approaches either operate directly in 3D, which ensures geometric consistency but often lacks generative expressiveness, or rely on 2D amodal completion, which provides strong appearance priors but does not guarantee reliable 3D structure. This raises a key question: how can we achieve both generative plausibility and geometric coherence in amodal 3D modeling? To answer this question, we introduce GENA3D (GENarative Amodal 3D), a framework that integrates learned 2D generative priors with explicit 3D geometric reasoning within a conditional 3D generation paradigm. The 2D priors enable the model to plausibly infer diverse occluded content, while the 3D representation enforces multi-view consistency and spatial validity. Our design incorporates a novel View-Wise Cross-Attention for multi-view alignment and a Stereo-Conditioned Cross-Attention to anchor generative predictions in 3D relationships. By combining generative imagination with structural constraints, GENA3D generates complete and coherent 3D objects from limited observations without sacrificing geometric fidelity. Experiments demonstrate that our method outperforms existing approaches in both synthetic and real-world amodal scenarios, highlighting the effectiveness of bridging 2D priors and 3D coherence in generating plausible and geometrically consistent 3D structures in complex environments.
Figure 1: Comparison of 3D object reconstruction/generation from sparse and partial occluded inputs. From left to right: input views, results from (a) MVS-based reconstruction (e.g., VGGT [wang2025vggt]), (b) 2D amodal completion lifted to 3D (TRELLIS [xiang2025structured]), and (c) GENA3D (ours). Our model faithfully generate complete, occlusion-free 3D geometry consistent with input partially visible constraints.
@@ 1 Introduction
Generating complete 3D objects under partial occlusions is a fundamental but challenging task in computer vision. In real-world scenarios, objects are often captured from arbitrary viewpoints, with certain portions hidden by surrounding structures or other objects. Such incomplete observations make the problem of generating 3D objects in the scene inherently amodal: a system must infer the geometry and appearance of regions that are never directly observed in the input images. This ability is crucial for object-level 3D scene understanding and downstream applications that are critically dependent on such capability, such as robotic interaction, AR/VR content creation, and embodied AI perception.
Recent advances in 3D object generation have been driven by diffusion-based and autoregressive models that learn powerful structural and semantic priors from large-scale 3D datasets. These approaches can synthesize diverse and realistic shapes from text prompts, single images, or even sparse inputs. However, most existing 3D generative models heavily depend on learned strong geometric priors and implicitly assume high-quality observations and canonical object poses to produce high-quality results. When confronted with sparse and partially occluded real-world inputs, these models often struggle to effectively leverage the limited visible evidence to plausibly infer the missing regions. The inherent complexity of the 3D domain and the scarcity of corresponding training data further exacerbate this issue, leading to inconsistent geometry, implausible structures, loss of fine-grained details, or even catastrophic generation failures.
@@
To address this challenge, we introduce GENA3D, a generative framework for amodal 3D object generation from sparse, unposed views. Our approach treats this problem as a conditional generation process that bridges 2D priors and 3D coherence. Specifically, we first perform 2D object-centric amodal completion that are rich in generative ability on each view to recover occluded regions. The features of completed but inconsistent views are then aggregated through a View-Wise Cross Attention module that produces model latents fused from sparse-view information. To ensure the fidelity of visible regions and 3D coherence, we condition the generation on partial geometry (point clouds) estimated from an MVS model. This information is incorporated through a Stereo-Conditioned Cross Attention module equipped with a geometry-guided gating mechanism, which modulates fused latent based on 3D spatial cues. After processing the generated sparse structure through another pretrained Flow Transformer, the resulting structured latent representation is then decoded to produce a complete, occlusion-free 3D model that faithfully respects to the observed geometry while plausibly generating the hidden regions. As shown in Fig. 1 (c), our method produces coherent and complete 3D objects even when large portions of the geometry are never observed or the 2D completions are not consistent.
@@
Extensive experiments on both synthetic and real-world datasets demonstrate that GENA3D achieves superior generation quality compared with existing sparse-view and amodal 3D generation baselines. Our method improves geometric faithfulness while maintaining realistic appearance, achieving consistent and superior performance from single-view to multi-view inputs. These results highlight the potential of generative modeling for practical 3D object-level scene synthesis under real-world occlusion and view sparsity.
@@
  • We introduce GENA3D, a generative amodal 3D modeling framework that recover complete geometry and appearance from sparse and occluded inputs by bridging 2D priors and 3D coherence during generation.
  • We propose two principled conditioning mechanisms for amodal 3D generation: (1) View-Wise Cross Attention, which introduces multi-view reasoning to mitigate view dominance and geometric drift; and (2) Stereo-Conditioned Cross Attention, which integrates partial stereo geometry through logit-level gating, transforming geometric cues into explicit attention regulators for structurally coherent generation.
  • Comprehensive experiments validate that GENA3D achieves superior generation fidelity, completeness, and consistency compared with existing sparse-view and amodal generation methods.
@@ 2 Related Work
3D Generative Models. Recent advances in 3D generative models have opened new possibilities for structure-aware reconstruction. Methods leveraging Gaussian Splatting [kerbl20233d] and NeRF [mildenhall2021nerf] as these representations excel at novel view synthesis [ye2024no, truong2023sparf], but struggle to segregate object-level entities, hindering object-level 3D scene understanding. Diffusion-based generative models further enhance visual fidelity, controllability, and scalability by leveraging pretrained 2D diffusion priors [huang2024dreamcontrol, ding2024bidirectional, chen2024vp3d] or 3D-aware latent representations [xiang2025structured, long2024wonder3d, zhao2025hunyuan3d, lin2025kiss3dgen, siddiqui2024assetgen]. These methods typically rely on massive optimization iterations or decently posed input images for proper results. Some works attempt to mitigate missing regions using inpainting-augmented generation [chen2024comboverse, han2024reparo, zhou2024layout, zhou2025coco4dcomprehensivecomplex4d]. Other scene-level generative pipelines [chen2024single, liu2022towards, huang2025midi, meng2025scenegen, yao2025cast] synthesize entire 3D scenes from a single image, but they rely on carefully curated data and pose-calibrated views, which limits their effectiveness under real-world, unposed scenarios.
@@
In contrast to prior approaches, GENA3D is a generative 3D reconstruction model tailored for sparse, unposed inputs. By fusing generative priors with partial geometric conditioning, it produces reconstructions that are both geometrically faithful and semantically coherent, remaining consistent with sparse-view constraints while plausibly inferring unseen object regions.
@@
Amodal Completion in 2D/3D. Amodal completion aims to infer the hidden or occluded parts of objects beyond visible evidence. Early works address this through explicit amodal segmentation and completion tasks [LiM_2016_ECCV, Qi_2019_CVPR, Ling_2020_NeurIPS, Mohan_2022_CVPR, Li_2023_ICCV_MUVA], focusing primarily on 2D perception and annotation-based benchmarks. These methods rely on supervised mask prediction or latent shape priors to hallucinate occluded object regions, but lack in fidelity and geometric consistency.
@@
Generative 2D inpainting models [Rombach_2022_CVPR, yu2023inpaint] have since become common solutions in image-level amodal completion, synthesizing realistic missing content based on surrounding context. Nowadays, researchers further enhance this capacity by introducing specialized amodal reasoning modules [ao2025open, lee2025tuningfreeamodalsegmentationocclusionfree, fan2025multi] or by fine-tuning diffusion models for occlusion-aware reconstruction [ozguroglu2024pix2gestalt, xu2024amodal, lu2025taco]. Extending beyond single-view imagery, multi-view inpainting models [barda2025instant3dit, weber2023nerfiller] attempt to maintain cross-view coherence but often assume calibrated cameras or dense view coverage. More geometry-aware approaches [Li_2022_ECCV, salimi2025geometry, feng2025objfiller, shi2025imfine, Zhan_2024_CVPR] leverage 3D representation priors for scene-level completion; however, they remain limited in handling object-level amodal generation due to alignment errors and weak semantic reasoning. The most relevant work is Amodal3R [wu2025amodal3r], which trains a native 3D generation model from occluded objects, but it is constrained by data scale and limited occlusion reasoning ability.
@@
In sparse and unposed scenarios, 2D inpainting models can recover locally plausible structures but often introduce cross-view inconsistencies that lead to geometric artifacts in 3D reconstruction. Our method integrates strong 2D amodal priors into a 3D generative framework, enforcing global geometric coherence via view-wise and stereo-conditioned attention. Compared with single-image 3D amodal methods, GENA3D enables more robust reconstruction under severe occlusion by aggregating multi-view information and enhancing cross-view geometric reasoning, while preserving realistic image-level completion.
Multi-View Stereo (MVS). Multi-view stereo (MVS) [FurukawaHernandez2015_MultiViewStereoTutorial, Wang2024_LearningBasedMVS_Survey, Yi2020PyramidMVSNet, Yan2020DenseHybridRMVSNet, wang2024dust3r, leroy2024grounding, zhang2024monst3r, yang2025fast3r, zhang2025flare, wang2025pi, keetha2025mapanything] has been studied for decades as a core technique for 3D reconstruction from multiple input views. Recent advances [Ren_2023_HPMMVS, Wu_2024_GoMVS, Santo_2024_MVCPS_NeuS] in dense correspondence learning have enhanced MVS accuracy and robustness by jointly reasoning about geometry and radiance. These methods achieve fine-grained alignment but remain limited when it comes to calibrated settings and fixed depth ranges. To improve scalability and avoid per-scene optimization, feed-forward extensions [Izquierdo_2025_MVSAnywhere, Cabon_2025_MUSt3R, Liu_2024_MVSGaussian, Dong_2024_DepthRangeFreeMVS] introduce transformer-based architectures and specific techniques for zero-shot multi-view reconstruction. These designs collectively enhance scalability and inference efficiency.
Despite substantial progress, MVS methods are primarily designed for reconstruction rather than generation. In contrast, we do not seek to improve MVS. Instead, we leverage partial MVS geometry as an explicit structural prior within a generative 3D framework. By conditioning on stereo-derived point clouds through our Stereo-Conditioned Cross Attention, GENA3D converts incomplete geometric cues into attention regulators that guide plausible and geometrically coherent completion. This effectively bridges feed-forward MVS geometry with conditional 3D generation for object-level amodal modeling.
Figure 2: Overview of GENA3D. Given sparse images $S$, visibility masks $M_{vis}$, and occlusion masks $M_{occ}$ indicating the occluded object $O$, GENA3D first generates a sparse structure (stage 1) by aggregating multi-view information in Sec. 3.1 and infers the complete geometric structure from the partial stereo point cloud $\mathbf{P}_{O}$ in Sec. 3.2. We then employ a pretrained amodal SLAT Transformer to generate details with $M_{vis}$ and $M_{occ}$, and then generate the structured latent. This structured latent is decoded into an occlusion-free 3D object with high-quality geometry and appearance.
@@ 3 Method
For an arbitrary object $O$ in the given $K$ sparse views $S=\{s^{1},s^{2},...,s^{K}\}$ (typically 1-4 views), we can obtain the visibility masks of $O$ and the corresponding text prompt automatically or manually using vision foundation models, e.g., SAM [kirillov2023segment] and Florence 2 [xiao2024florence]. After that, we follow the occlusion analysis and 2D amodal completion strategies from [ao2025open] with minor modifications (using the above obtained visibility mask and prompt) to first perform generative image-level amodal completion across all views using  [ao2025open] (denoted as OAAC). Altogether, we can obtain visual clues of the desired object $O$ at all $K$ input views: the visibility masks $M_{vis}=\{m^{1}_{v},m^{2}_{v},...,m^{K}_{v}\}$ indicating the visible part of $O$; the inconsistent 2D completions of $O$: $I=\{i^{1},i^{2},...,i^{K}\}$; and the occlusion masks $M_{occ}=\{m^{1}_{o},m^{2}_{o},...,m^{K}_{o}\}$ representing the potential occlusion regions.
We illustrate the overall model structure in Fig. 2. Our design prioritizes the Sparse Structure Transformer, i.e., stage 1. Given object-centric sparse views $I$ processed using 2D amodal completion methods, we apply the proposed View-Wise Cross Attention (Sec. 3.1) to aggregate multi-view features and fuse their latents. The fused representation is further refined through the Stereo-Conditioned Cross Attention (Sec. 3.2) for unobserved structure inference, guided by observed partial stereo geometry and a geometry-aware gating mechanism. The resulting Sparse Structure is decoded and then passed to the pretrained Structure Latent Transformer from [wu2025amodal3r] for Structured Latent generation. Implementation details and training data preparation are described in Sec. 3.3.
Figure 3: The original model shows a strong bias and results in artifact accumulation when conditioned with multi-view.
Figure 4: A detailed illustration of our proposed View-Wise Cross Attention and Stereo-Conditioned Cross Attention modules.